x/3=1/6 I know the answer is 1/2 but cannot explain the process to my son. Pls help with steps I am old!
Answer by stanbon(75887)

x/3=1/6
Multiply both sides of the equation by 3 to get:
x=3/6
Reduce the fraction to get:
x=1/2
Cheers,
Stan H.